# T-FIX: Text-Based Explanations with Features Interpretable to eXperts

### v1

@@ 5 Included Datasets
GPT-4o (openai2024gpt4ocard)
Vanilla 0.334 0.726 0.564 0.494 0.236 0.427 0.535
Chain-of-Thought 0.320 0.697 0.534 0.514 0.278 0.472 0.541
Socratic Prompting 0.329 0.708 0.536 0.501 0.299 0.466 0.547
SubQ Decomposition 0.279 0.719 0.523 0.439 0.294 0.435 0.553
o1 (openai2024openaio1card)
Vanilla 0.528 0.570 0.560 0.521 0.370 0.446 0.525
Chain-of-Thought 0.501 0.552 0.562 0.555 0.387 0.428 0.560
Socratic Prompting 0.428 0.575 0.502 0.397 0.371 0.370 0.555
SubQ Decomposition 0.430 0.582 0.482 0.480 0.328 0.398 0.574
Gemini-2.0-Flash (google2024gemini2flash)
Vanilla 0.454 0.650 0.552 0.495 0.335 0.445 0.544
Chain-of-Thought 0.442 0.666 0.505 0.489 0.312 0.465 0.539
Socratic Prompting 0.223 0.666 0.490 0.470 0.323 0.409 0.560
SubQ Decomposition 0.323 0.660 0.509 0.449 0.325 0.408 0.557
Claude-3.5-Sonnet (anthropic2025claude35)
Vanilla 0.622 0.581 0.558 0.533 0.211 0.459 0.564
Chain-of-Thought 0.605 0.604 0.565 0.493 0.228 0.440 0.563
Socratic Prompting 0.616 0.613 0.529 0.474 0.235 0.448 0.557
SubQ Decomposition 0.532 0.594 0.554 0.503 0.221 0.460 0.530
Table 3: Evaluating top LLMs on T-FIX. We report the average expert alignment score across all examples in the dataset. Corresponding accuracies are in Table˜A2 and baseline prompting strategies are described in Section 6. In general, more sophisticated prompting strategies do not necessarily improve expert alignment of explanation.
T-FIX contains seven open-source datasets, spanning the fields of cosmology, psychology, and medicine. To assess LLM explanations across multiple modalities, we include text, vision, and time-series datasets. We select these seven datasets due to the availability of domain experts willing to work with us for these tasks.
As running T-FIX requires querying LLMs, many of which follow a pay-as-you-go API structure, we cap the benchmark at 700 (100 per dataset) in order for T-FIX to be accessible to as many researchers as possible.
We select a subset of 100 examples from the test set of each open-source dataset in T-FIX, and balance this sampling across classes when possible. We provide an overview of the included open-source datasets in Figure 4. See Appendix˜D for additional details about the motivation, task, and prompting procedure for each dataset.
@@ 6 Experiments
After building a pipeline to evaluate the expert alignment of an LLM explanation, we evaluate a suite of today’s top LLMs on T-FIX to determine how expert-aligned these models are on domain-specific tasks. We use the following prompting techniques as baselines to generate explanations for each dataset in T-FIX.
@@
    Chain-of-Thought (CoT): The LLM is prompted to reason step-by-step through intermediate steps before answering, supporting more accurate responses on multi-step tasks.
@@
    Subquestion Decomposition: The LLM is guided to break down a complex task into simpler subquestions, answer them individually, and then synthesize a final response.
Domain-specific prompts are in Appendix˜D, with templates for the above prompting strategies in Figure˜A5. Results for GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet  [^3] We only select LLMs with vision support and context windows long enough to accommodate our time-series datasets. All models are accessed in September 2025. are in Table˜3.
Role of LLM in pipeline.
Because our prompts are tightly constrained and refined through multiple iterations, the evaluation pipeline is largely robust to the choice of LLM we use. We validate this by comparing results across three evaluator models — GPT-4o, Gemini-2.0-Flash, and Qwen2.5-VL-7B-Instruct — on one dataset per domain. Despite minor variations, rankings and alignment score ranges remain consistent, confirming that the pipeline is robust to evaluator choice. Further details are provided in Appendix˜C.
@@ 7 Analysis
In this section, we analyze how LLMs distribute reasoning across expert criteria and whether higher task accuracy indicates better expert alignment.
Figure 5: Shannon Entropy of expert alignment criteria for GPT-4o. For each prompting baseline, we show coverage of each domain’s explanations across all expert criteria – a high value indicates the LLM considers many criteria across examples, while a low value indicates the LLM focuses on the same criteria repeatedly.
Figure 6: Expert Alignment vs. Accuracy Correlation Heatmap, averaged across GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet. Blue indicates positive correlation, red is negative, gray is no correlation.
@@ 7.1 Coverage of Expert Alignment Criteria
Figure 5 presents the Shannon entropy of GPT-4o’s covered expert criteria in each domain. We observe a correlation between entropy and performance: domains where GPT-4o underperforms (e.g., Cholecystectomy, Supernova) show lower entropy, indicating limited criteria coverage. In contrast, well-performing domains (e.g., Politeness, Sepsis) exhibit more uniform coverage, equally taking into account all expert criteria.
This suggests that LLMs that reason uniformly over expert alignment criteria perform better – a promising insight for future work in prompting or training models to incorporate a broader range of expert reasoning.
@@
T-FIX focuses on evaluating explanation quality, but we are also interested in understanding the relationship between expert alignment and prediction accuracy. Specifically, we ask: Does higher answer accuracy correspond to stronger expert alignment?
Figure 6 shows the Pearson correlation of expert alignment (see Table˜3) with accuracy (see Table˜A2) for each domain, averaged across models. In some domains with higher performance, like Cholecystectomy and Emotion, we do observe higher expert alignment as well. However, the overall correlation is weak across domains.
The heatmap suggests today’s high-performing LLMs do not appear to rely on expert reasoning. Future research is needed to explore whether aligning model reasoning with expert criteria – via training objectives or prompting – can improve downstream performance.
@@ 8 Related Work
Evaluating LLM explanations.
Common explanation methods for LLMs include feature attribution (e.g., LIME, SHAP ribeiro2016lime; lundberg2017shap), counterfactuals, and self-generated explanations im2023evaluatingutilitymodelexplanations; zhao2023explainabilitylargelanguagemodels. Some models are also trained to produce human-readable justifications camburu2018esnli. To assess explanation quality and utility, recent work highlights criteria such as faithfulness (alignment with the model’s reasoning) and plausibility (how convincing it is to humans) jacovi2020faithfulness; zhou2021evaluating; agarwal2024faithfulness. Human studies show mixed outcomes: explanations sometimes aid understanding hase2020evaluating; bansal2021whole, but can also offer little value or cause over-trust wang2023fair. A promising alternative is to use LLMs as automatic judges of explanation quality zheng2023judging; chen2024judge, providing a scalable substitute for expensive human evaluation; we adopt this approach in T-FIX.
Domain & expert alignment.
Concept-based models constrain parts of the network to predict high-level, human-defined concepts, enabling incorporation of domain knowledge into final predictions koh2020concept. Extensions of concept bottlenecks and related methods aim to align latent representations with semantically meaningful features kim2018tcav; chen2020concept; ghorbani2019ace, potentially grouped for expert interpretability jin2024fix. In NLP, integrating human knowledge has included collecting human-written explanation datasets to train models camburu2018esnli and using learned explanations to guide predictions bhatt2020explainable. To our knowledge, no prior work explicitly evaluates text explanations for expert alignment like T-FIX.
@@ 9 Conclusion
We introduce T-FIX, the first benchmark designed to evaluate LLM explanations for expert alignment across seven knowledge-intensive domains. Our analysis reveals that today’s models struggle to generate explanations that experts would rely on, highlighting a critical area for improvement.
Future work may include exploring instruction-tuning LLMs to generate explanations with strong expert alignment, extending T-FIX to additional domains, and Human-Computer Interaction studies exploring how expert-aligned explanations affect real-world decision-making by practitioners.
@@
While our evaluation pipeline currently uses GPT-4o for scoring, it is model-agnostic by design, and we encourage future work to apply or adapt the pipeline with other LLMs to improve robustness and reduce evaluator-model entanglement.
@@
For pipeline validation, we conduct a user study where we annotate 35 examples. Though the annotation results on this subset suggest our pipeline is accurate, this work could have benefited from a larger and more robust annotation study. Future work should also involve domain experts vetting the pipeline in addition to recruited annotators.
In addition, we only have one expert to validate the expert alignment criteria for each domain. Although our usage of a deep research LLM minimizes over-reliance on a single domain expert, multiple experts would have been better to create the expert criteria. We were constrained by the availability of domain experts.
@@
Our experiments focus on a set of four models and four prompting strategies, and including additional models and strategies could provide a more comprehensive set of baseline results. Though many other high-performing LLMs and prompting techniques exist as of September 2025, we are conscious of budget and the environmental impact of running multiple experiments using T-FIX.
Ethical Considerations
@@
Using LLMs in the domains we describe in T-FIX, especially those relating to medicine, poses a unique set of risks and challenges. We do not advocate that LLMs should replace domain experts in these tasks; rather, T-FIX should serve as a step towards experts being able to use LLMs in a reliable and trustworthy way.
Additionally, LLMs are constantly changing, especially those that are company-owned and not open-source. This poses potential issues relating to the reproducibility of our baseline results as time progresses and advances are made.
Lastly, nearly all LLMs contain biases – some harmful – that may propagate up in a system built off of these models. All users of T-FIX must be conscious of this risk.
Evaluation Model
GPT-4o
Gemini-2.0-Flash
Qwen2.5-VL-7B-Instruct
(Qwen2.5-VL)
Mass Maps
GPT-4o 0.334 0.394 0.507
o1 0.528 0.581 0.596
Gemini-2.0-Flash 0.454 0.483 0.431
Claude-3.5-Sonnet 0.622 0.692 0.651
Emotion
GPT-4o 0.494 0.595 0.579
o1 0.521 0.632 0.567
Gemini-2.0-Flash 0.495 0.621 0.567
Claude-3.5-Sonnet 0.533 0.610 0.565
Cholecystectomy
GPT-4o 0.236 0.000 0.145
o1 0.370 0.000 0.234
Gemini-2.0-Flash 0.335 0.000 0.153
Claude-3.5-Sonnet 0.211 0.000 0.146
Table A1: Model ablation study for evaluation pipeline. Across domains, model rankings are largely consistent across evaluators, indicating the robustness of our evaluation pipeline. Comparison of base models (explanation generator LLMs) across evaluator models (pipeline LLMs) for one dataset per domain. Best per evaluator is in bold; second-best is underlined.

### v2

@@ 5 Included Datasets
Table 2: Evaluating top LLMs on T-FIX using GPT-5-mini. We report the average $\mathsf{T\textrm{-}FIXscore}$ across all examples in the dataset. Corresponding accuracies are in Table˜A2 and baseline prompting strategies are described in Section 6. We report $\pm$ standard deviation over 1000 bootstrap samples. In general, more sophisticated prompting strategies do not improve expert alignment of the explanations.
@@
GPT-5.2-Pro (OpenAI, 2025a)
Vanilla $0.901_{\pm 0.01}$ $0.845_{\pm 0.01}$ $0.882_{\pm 0.01}$ $0.906_{\pm 0.01}$ $0.538_{\pm 0.03}$ $0.539_{\pm 0.02}$ $0.425_{\pm 0.03}$
[…]
Socratic Prompting $0.804_{\pm 0.01}$ $0.714_{\pm 0.01}$ $0.763_{\pm 0.02}$ $0.726_{\pm 0.02}$ $0.656_{\pm 0.02}$ $0.488_{\pm 0.02}$ $0.381_{\pm 0.02}$
SubQ Decomposition $0.812_{\pm 0.01}$ $0.727_{\pm 0.01}$ $0.857_{\pm 0.01}$ $0.774_{\pm 0.02}$ $0.639_{\pm 0.02}$ $0.489_{\pm 0.02}$ $0.356_{\pm 0.02}$
GPT-5-mini (OpenAI, 2025b)
[…]
Chain-of-Thought $0.759_{\pm 0.01}$ $0.876_{\pm 0.01}$ $0.767_{\pm 0.02}$ $0.830_{\pm 0.02}$ $0.609_{\pm 0.03}$ $0.507_{\pm 0.02}$ $0.395_{\pm 0.03}$
Socratic Prompting $0.729_{\pm 0.01}$ $0.781_{\pm 0.01}$ $0.641_{\pm 0.02}$ $0.690_{\pm 0.02}$ $0.506_{\pm 0.03}$ $0.444_{\pm 0.02}$ $0.399_{\pm 0.02}$
SubQ Decomposition $0.720_{\pm 0.02}$ $0.838_{\pm 0.01}$ $0.809_{\pm 0.02}$ $0.846_{\pm 0.02}$ $0.528_{\pm 0.03}$ $0.443_{\pm 0.01}$ $0.419_{\pm 0.03}$
Claude-Opus-4.5 (Anthropic, 2025b)
[…]
Socratic Prompting $0.835_{\pm 0.01}$ $0.683_{\pm 0.01}$ $0.802_{\pm 0.02}$ $0.789_{\pm 0.02}$ $0.647_{\pm 0.01}$ $0.653_{\pm 0.02}$ $0.390_{\pm 0.02}$
SubQ Decomposition $0.765_{\pm 0.01}$ $0.736_{\pm 0.01}$ $0.800_{\pm 0.02}$ $0.821_{\pm 0.02}$ $0.546_{\pm 0.01}$ $0.620_{\pm 0.02}$ $0.386_{\pm 0.02}$
Claude-Haiku-4.5 (Anthropic, 2025a)
Vanilla $0.742_{\pm 0.01}$ $0.630_{\pm 0.02}$ $0.873_{\pm 0.02}$ $0.881_{\pm 0.02}$ $0.677_{\pm 0.01}$ $0.529_{\pm 0.02}$ $0.355_{\pm 0.02}$
Chain-of-Thought $0.781_{\pm 0.01}$ $0.640_{\pm 0.02}$ $0.871_{\pm 0.02}$ $0.887_{\pm 0.02}$ $0.722_{\pm 0.01}$ $0.510_{\pm 0.02}$ $0.368_{\pm 0.02}$
Socratic Prompting $0.686_{\pm 0.01}$ $0.658_{\pm 0.02}$ $0.838_{\pm 0.01}$ $0.842_{\pm 0.01}$ $0.687_{\pm 0.01}$ $0.492_{\pm 0.02}$ $0.300_{\pm 0.02}$
SubQ Decomposition $0.626_{\pm 0.02}$ $0.537_{\pm 0.02}$ $0.769_{\pm 0.02}$ $0.841_{\pm 0.02}$ $0.616_{\pm 0.01}$ $0.505_{\pm 0.02}$ $0.308_{\pm 0.02}$
Gemini-2.5-Pro (team, 2025)
Vanilla $0.801_{\pm 0.01}$ $0.715_{\pm 0.02}$ $0.843_{\pm 0.02}$ $0.888_{\pm 0.02}$ $0.697_{\pm 0.01}$ $0.315_{\pm 0.02}$ $0.511_{\pm 0.02}$
Chain-of-Thought $0.812_{\pm 0.01}$ $0.697_{\pm 0.02}$ $0.877_{\pm 0.02}$ $0.873_{\pm 0.02}$ $0.691_{\pm 0.02}$ $0.264_{\pm 0.02}$ $0.525_{\pm 0.02}$
Socratic Prompting $0.671_{\pm 0.01}$ $0.616_{\pm 0.02}$ $0.682_{\pm 0.02}$ $0.655_{\pm 0.02}$ $0.670_{\pm 0.01}$ $0.281_{\pm 0.02}$ $0.466_{\pm 0.02}$
SubQ Decomposition $0.731_{\pm 0.01}$ $0.635_{\pm 0.02}$ $0.759_{\pm 0.02}$ $0.836_{\pm 0.02}$ $0.654_{\pm 0.01}$ $0.221_{\pm 0.01}$ $0.522_{\pm 0.02}$
Gemini-2.5-Flash (team, 2025)
[…]
Socratic Prompting $0.790_{\pm 0.01}$ $0.683_{\pm 0.02}$ $0.831_{\pm 0.02}$ $0.812_{\pm 0.02}$ $0.620_{\pm 0.01}$ $0.315_{\pm 0.02}$ $0.484_{\pm 0.02}$
SubQ Decomposition $0.800_{\pm 0.01}$ $0.688_{\pm 0.02}$ $0.816_{\pm 0.02}$ $0.841_{\pm 0.02}$ $0.688_{\pm 0.01}$ $0.300_{\pm 0.02}$ $0.506_{\pm 0.02}$
T-FIX aggregates seven open-source datasets, spanning the fields of cosmology, psychology, and medicine. To assess LLM explanations across multiple modalities, we include text, vision, and time-series datasets. We select these seven datasets due to the availability of domain experts willing to work with us for these tasks.
As running T-FIX requires querying LLMs, many of which follow a pay-as-you-go API structure, we cap the benchmark at 700 (100 per dataset) in order to manage costs and enable T-FIX to be accessible to as many researchers as possible. We select 100 examples from the test set of each open-source dataset in T-FIX, and balance this sampling across classes when possible. We provide an overview of the included open-source datasets in Figure 4. See Appendix˜G for additional details about the motivation, task, and prompting procedure for each dataset.
@@ 6 Experiments
After building a pipeline to evaluate the expert alignment of an LLM explanation, we evaluate a suite of today’s top LLMs on T-FIX to determine how expert-aligned these models are on domain-specific tasks.
We use the following prompting techniques as baselines to generate explanations for each dataset in T-FIX.
@@
    Chain-of-Thought (CoT): The LLM is prompted to reason step-by-step through intermediate steps before answering, improving responses on multi-step tasks.
@@
    Subquestion Decomposition: The LLM is guided to break down the task into simpler subquestions, answer them individually, and then synthesize a final response.
Domain-specific prompts are in Appendix˜G, with templates for the above prompting strategies in Figure˜A6. Results for GPT-5.2, GPT-5-mini, Gemini-2.5-Flash, Gemini-2.5-Pro, Claude-Opus-4.5, and Claude-Haiku-4.5222We only select stable LLMs (i.e., not in preview mode) with vision support and context windows long enough to accommodate our time-series datasets. All models are accessed in January 2026. are in Table˜2.
Role of evaluator LLM. Because our prompts are tightly constrained and refined through multiple iterations, the evaluation pipeline is largely robust to the choice of LLM we use. We validate this by comparing results across three evaluator models: GPT-5-mini, Gemini-2.5-Flash-Lite, and Qwen2.5-VL-7B-Instruct, on one dataset per domain.
Despite minor variations, rankings and alignment score ranges remain consistent, confirming that the pipeline is robust to evaluator choice. Further details are provided in Appendix˜C.
@@ 7 Analysis
In this section, we analyze how LLMs distribute reasoning across expert criteria and whether higher expert alignment indicates higher accuracy.
Figure 5: Shannon Entropy of expert alignment criteria for GPT-5-mini. For each prompting baseline, we show coverage of each domain’s explanations across all expert criteria – a high value indicates the LLM considers many criteria across examples, while a low value indicates the LLM focuses on the same criteria repeatedly.
Figure 6: Expert Alignment vs. Accuracy Correlation Heatmap, averaged across all 6 models in Table˜2. Blue indicates positive correlation, red is negative, gray is no correlation. $^{*}p<0.05$.
@@ 7.1 Coverage of Expert Alignment Criteria
Figure 5 presents the Shannon entropy of GPT-5-mini’s covered expert criteria in each domain. We observe a correlation between entropy and performance: domains where GPT-5-mini underperforms (e.g., Cholecystectomy, Supernova) show lower entropy, indicating limited criteria coverage. In contrast, higher-performing domains (e.g., Politeness, Sepsis) exhibit more uniform coverage, equally taking into account all expert criteria.
This suggests that LLMs that reason uniformly over expert alignment criteria perform better — a promising insight for future work in prompting or training models to incorporate a broader range of expert reasoning.
@@
T-FIX evaluates explanation quality, but we also analyze how expert alignment relates to prediction accuracy. Specifically, we ask: Does higher answer accuracy correspond to stronger expert alignment?
Figure 6 shows Pearson correlations between expert alignment (Table˜2) and accuracy (Table˜A2), averaged across models. Across domains, the vast majority of correlations are statistically insignificant or low in magnitude.
The heatmap suggests today’s high-performing LLMs do not appear to rely on expert reasoning. Future research is needed to explore whether explicitly aligning model reasoning with expert criteria — via training objectives or prompting — can improve downstream performance.
@@ 8 Related Work
Evaluating LLM explanations at scale. Common methods include self-generated explanations (Im et al., 2023; Zhao et al., 2023) or human-readable justifications (Camburu et al., 2018). To assess explanation quality and utility, faithfulness is analyzed (Jacovi and Goldberg, 2020; Zhou et al., 2021; Agarwal et al., 2024). Human studies show mixed outcomes: explanations sometimes aid understanding (Hase and Bansal, 2020; Bansal et al., 2021), but can also offer little value or cause over-trust (Wang et al., 2023). A promising alternative is to use LLMs as automatic judges of explanation quality (Zheng et al., 2023; Chen et al., 2024), providing a scalable substitute for expensive human evaluation; we adopt this approach in T-FIX.
Domain & expert alignment. Prior work incorporates domain knowledge through concept-based representations that constrain or structure predictions (Koh et al., 2020; Kim et al., 2018; Chen et al., 2020; Ghorbani et al., 2019; Jin et al., 2025), and through explanation supervision in NLP (Camburu et al., 2018; Bhatt et al., 2020). Existing efforts to evaluate expert alignment are intensive and small-scale (Kutbi et al., 2025; Hyk et al., 2025)), while larger domain benchmarks rarely assess expert alignment beyond validation (e.g., Astro-QA (Li et al., 2025a), CounselBench (Li et al., 2025b), Mentalchat16k (Xu et al., 2025)) and rubric-based medical evaluations remain tied to fixed examples (Yang et al., 2026). To our knowledge, no prior work explicitly evaluates expert alignment like T-FIX does.
@@ 9 Conclusion
We introduce T-FIX, the first benchmark designed to evaluate LLM explanations for expert alignment across seven knowledge-intensive domains. Unlike existing evaluation frameworks, T-FIX front-loads expert effort into a reusable framework that enables automatic expert-alignment scoring of unseen explanations. Future work may include exploring instruction-tuning LLMs to generate explanations with strong expert alignment, extending T-FIX to additional domains, and Human-Computer Interaction studies exploring how expert-aligned explanations affect real-world decision-making by practitioners.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
While our evaluation pipeline currently uses GPT-5-mini for scoring, it is model-agnostic by design, and we encourage future work to apply or adapt the pipeline with other LLMs to improve robustness and reduce evaluator-model entanglement. For pipeline validation, we conduct a user study where we annotate 35 examples. Though the annotation results on this subset suggest our pipeline is accurate, this work could have benefited from a larger and more robust annotation study. Future work should also involve domain experts vetting the pipeline in addition to recruited annotators. In addition, we only have a small number of experts to validate the expert alignment criteria for each domain. Having additional experts would have been better to create a more robust expert criteria. We were constrained by the availability of domain experts.
@@
Our experiments focus on a set of six models and four prompting strategies, and including additional models and strategies could provide a more comprehensive set of baseline results. Though many other high-performing LLMs and prompting techniques exist as of January 2026, we are conscious of budget and the environmental impact of running multiple experiments using T-FIX. Additionally, LLMs are constantly changing, especially those that are company-owned and not open-source. This poses potential issues relating to the reproducibility of our baseline results as time progresses and advances are made.
@@
Finally, using LLMs in the domains we describe in T-FIX, especially those relating to medicine, poses a unique set of risks and challenges. We do not advocate that LLMs should replace domain experts in these tasks; rather, T-FIX should serve as a step towards experts being able to use LLMs in a reliable and trustworthy way.
Appendix A Extending T-FIX to a New Domain
To add a new domain to T-FIX, we advise you to follow these steps:
  1. 1.
    Generate criteria: Use the deep research prompt template shown in Figure˜A5 to generate a list of expert alignment criteria for your domain. Optionally, have a domain expert vet the generated criteria.
  2. 2.
    Modify prompts: Modify the prompt templates outlined in Figure˜A1, Figure˜A2, Figure˜A3, and Figure˜A4 with your task description, few-shot examples, and generated expert criteria.
  3. 3.
Appendix B Annotation Study Setup
Annotator qualifications. Annotators are PhD students in machine learning and the authors of this paper. Each author annotating for a scientific domain has had multiple meetings with that domain’s experts, and been taught exactly how to evaluate an LLM’s explanation for expert alignment. Note that this annotation study is meant to support the output of the T-FIX pipeline in conjunction with the qualitative study described in Appendix D.
Annotators evaluate each stage using a unified labeling scheme: (A) fully accurate, (B) partially accurate, or (C) incorrect.
For atomic claim extraction, they review the explanation and its extracted claims, labeling outputs as (A) correct decomposition, (B) 1–2 claims missing/incorrect, or (C) 3+ claims missing/incorrect.
For filtering, they assess whether claims were appropriately kept or removed, labeling results as (A) all correct, (B) 1–2 incorrect, or (C) 3+ incorrect.
For expert alignment scoring, annotators review the grouped claims and expert criteria, labeling each as (A) consistent with expert-defined criteria, (B) borderline, or (C) inconsistent with expert-defined criteria.
These labels are then mapped to accuracy scores of 1.0, 0.5, and 0.0, respectively. Table˜A1 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX datasets, indicating moderate-to-substantial agreement and supporting the reliability of our pipeline. Domain-level metrics appear in Table˜A3.
Table A1: Pipeline validation: Average accuracy for $\mathcal{N}$ samples across all T-FIX domains and annotator agreement – Cohen’s $\kappa$ for each stage in our pipeline. Domain-specific statistics are provided in Table˜A3.
Pipeline Stage $\mathcal{N}$ Accuracy Cohen’s $\kappa$
Claim Extraction 35 0.943 0.717
Filtering & Grouping 295 0.871 0.402
Expert Alignment 211 0.923 0.405
Table A2: Evaluating top LLMs on T-FIX. We report the average performance of the LLM across all examples in the dataset. We report accuracy for classification tasks, and MSE for regression tasks – a () indicates that the score reported is MSE. Baseline implementations are described in Section 6. We report $\pm$ standard deviation over 1000 bootstrap samples.